OpenStudy (anonymous):

The width of a rectangle is 1 ft less than the length. The area is 2 ft^2. The length is _ The width is _ is the Length 1 and width 2?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

length= L width =L-1 L^2-L =2 L^2-L-2 =0 (L-2)(L+1) L=2 since ther is no negative length length =2 width =1

OpenStudy (lalaly):

your wrong... length is 2 and width is 1

OpenStudy (anonymous):

K. Thanks